OpenStudy (narissa):

help solve the equation 4x^2-100=0

OpenStudy (narissa):

@Directrix

OpenStudy (kayders1997):

First what can you do to this equation because it is set equal to 0

OpenStudy (narissa):

inverse operation

OpenStudy (kayders1997):

You don't need to make it to complicated, you can factor though

OpenStudy (narissa):

im confused

OpenStudy (kayders1997):

The greatest common factor

OpenStudy (kayders1997):

What is something that is common between 4x^2 and 100

OpenStudy (narissa):

idk 6

OpenStudy (narissa):

@umerlodhi

umerlodhi (umerlodhi):

one sec i gochu

umerlodhi (umerlodhi):

x = +5

OpenStudy (narissa):

thats what i thought thanks @umerlodhi

umerlodhi (umerlodhi):

1 Add 100100 to both sides 4x2=100 4x2=100 2 Divide both sides by 44 x2=1004 x2=1004 3 Simplify 10041004 to 2525 x2=25 x2=25 4 Take the square root of both sides x=±25−−√ x=±25 5 Since 5×5=255×5=25, the square root of 2525 is 55 x=±5

umerlodhi (umerlodhi):

ur welcome medal pls

OpenStudy (anonymous):

\[4x ^{2} - 100 <=> x^{2} - 400\] factor : \[4(x-5)^{2}\] cmiiw

umerlodhi (umerlodhi):

ty new thread pls

OpenStudy (anonymous):

so \[x = \pm5\]